OpenStudy (anonymous):

Exponents and logarithms and radical numbers...sooo simplify the expression nr.4

OpenStudy (anonymous):

1)\[\frac{ 1 }{2-\sqrt{3} }\] 2)\[\frac{ 1 }{ 3+\sqrt{5} }\]

OpenStudy (anonymous):

You need to find the conjugate and multiply the top and bottom by it.

OpenStudy (anonymous):

The conjugate of \(a+\sqrt{b} \) is \(a-\sqrt{b}\)

OpenStudy (anonymous):

and vice versa.

OpenStudy (anonymous):

Can you do this for 1?

OpenStudy (anonymous):

@ojeano I need cooperation here.

OpenStudy (anonymous):

aby rat5ionalizing we get, \[[2+\sqrt{3}]/[(2)^2-(\sqrt{3}^2)]\]

OpenStudy (anonymous):

(2+root3)/(4-3) =(2+root3)

OpenStudy (anonymous):

similarly try to find out second one.